Let Xi iid∼ Unif(0, θ) for i = 1, 2, 3 and

theta hat = 4Y(1)

theta tilde = 2Y(2)

θ* = (4/3)Y(3)

(a) For each estimator compute the bias?

(b) For each estimator compute the variance?

(c) Which estimator is best and why?
Answer #1

Define y Xcny mae {Xy,.Xn}, OS Xcmg<0. [: xicity (0,0)] . cids of Xay: Fix a ) - P(Xeg&2) = P(max {x{} <x) •(PCX,$2)) [:: moDefine, Key = min {xy , xn, 0x Koyko [= xi tiduros] -eds of Xco: Fix com = P(Meycz) = 1P(4,372)= | 4P(*1x2) Li min & Xı, -Xn53 - 2 a $6.21 022 p pdf of order statistic : X (2) if X,,..,Xn are in samples. e pucaseros hora CE 2) ** (1-7(a)ma face) [F(aSO ECC*)- E (VG)) = (*)=429.50. [Note: E(az)-a E(2)]. -: Bino (o)= ECO-0 =0 s) var( ) = Vaz (4 Ycw) = 4* Van (kent) = 16.92